OpenStudy (anonymous):

Please help: Find an explicit rule for the nth term of a geometric sequence where the second and fifth terms are 40 and 5000, respectively.

OpenStudy (anonymous):

an = 8 • 5^n - 2 an = 8 • 5^n + 3 an = 8 • 5^n - 1 an = 8 • 5^n + 1 I think it's this one

OpenStudy (anonymous):

@cwrw238